# Towards Fairness-Aware Adversarial Learning

### v1

@@ 4.2 Training from Scratch for Enhancing Fairness
Previous sections demonstrate the effectiveness and efficiency of the proposed approaches. Here we also investigate the advancements by training the model from the ground up using our method. We compare our methods with two common adversarial training methods (PGD-AT [28] and TRADES [44]) and two recent state-of-the-art techniques: CFA [39] and WAT [25], which have been proposed to mitigate the robust fairness issues recently. We adversarially trained Preact-ResNet-18 models [16] for 200 epochs with a learning rate of 0.1, which will be decayed by a factor of 0.1 at 100 and 150 epochs, successively. We start to facilitate the proposed intermediate maximization (see Algorithm 1 lines 9-14) after the 100-th epoch with the only hyper-parameter from 0.25 and enlarge it to 0.5 after the 150-th epoch. In addition, similar to CFA using weight average, we also applied EMA [20, 37], to gain a more stable performance, however, we only applied it after the 100-th epoch, where we start to apply the intermediate maximization. We report the best results under AutoAttack on the average accuracy and worst-class accuracy in Tab. 2. Besides, Fig. 3 visualizes the results of different training approaches including the proposed FAAL with other 3 TRADES-based models: TRADES, CFA and WAT respectively. We can observe that FAAL outperforms other approaches on the worst-class clean/robust accuracy, with less sacrifice on the average robustness.
@@
| + Fine-tune with FRL-RWRM | 52.55 (22.00) | 21.11 (2.00) |
| + Fine-tune with FAAL | 58.50 (21.00) | 21.91 (2.00) |
| + Fine-tune with FAAL | 58.41 (19.00) | 23.44 (2.00) |
| + Fine-tune with FAAL | 54.96 (18.00) | 22.71 (2.00) |
| + Fine-tune with FAAL | 54.90 (18.00) | 23.25 (2.00) |
| CFA | 55.57 (23.00) | 24.56 (2.00) |
| WAT | 53.99 (19.00) | 22.89 (3.00) |
| FAAL | 56.84 (16.00) | 21.85 (3.00) |
| FAAL | 55.87 (21.00) | 23.57 (3.00) |
@@ 4.3 Additional Experiments on CIFAR-100 dataset
The experiments above mainly focused on CIFAR-10 dataset, which only has 10 classes in the dataset. In this section, we explore the proposed FAAL into a more challenging dataset, i.e. CIFAR-100 with 100 categories. Similarly, we reported the results of the average/worst clean accuracy and AutoAttack accuracy. The value of in our method is set to 0.05. For fine-tuning, we compare our proposed method FAAL with FRL-RWRM [42], it can be seen that FAAL is able to achieve comparable to FRL-RWRM while reducing the amount of learning epoch up to 40 times (2 epochs vs. 80 epochs). For full adversarial training, following the experimental settings in WAT [25], we train the ResNet-18 models for 100 epochs via different adversarial training approaches, where the learning rate is decayed from 0.1 to 0.01 and 0.001 at the 75-th epoch and the 90-th epoch, respectively. We compare the results of FAAL compared to three baselines, i.e. TRADES, CFA and WAT. It can be seen that FAAT achieves the highest worst-class robust accuracy (same as WAT), meanwhile, it remains comparable results on the average robustness without sacrificing the average/worst-class clean accuracy. More details of the training settings can be found in the Appendix.
@@ 5 Essential Differences to SOTAs
Unlike those methods that rely on historical performance or an extra validation set for manual or heuristic weight adjustment per class in each epoch, our method bypasses these requirements. FAAL introduces an additional conic convex optimization problem after the adversarial example generation, based solely on the current batch’s objective loss, the bringing solving cost is negligible. The comparison of training computation time and other key properties is illustrated in Tab. 4. As model training can be unpredictable due to random mini-batch sampling, causing quick shifts in class distribution and bias that may differ from previous epochs or validations More importantly, FAAL can generalize to any adversarial training methods, as our intermediate maximization is a completely dependent component of the popular min-max framework, so it is not limited to any adversaries, unlike some methods FRL and WAT that are restricted to TRADES. Our data-driven component enhances flexibility in managing the balance between average robustness and robust fairness during adversarial training, and demonstrates its potential in handling various distribution shifts for the current batch.
@@ 6 Conclusion
In conclusion, we establish a connection between robust fairness and group distribution shift overfitting and present a new fairness-aware adversarial learning paradigm to address robust fairness. Compared to state-of-the-art methods, extensive experiments on CIFAR-10 and CIFAR-100 datasets demonstrate the effectiveness and superior efficiency of the proposed approach. Notably, by just two epochs of fine-tuning, our training strategy can transform a biased robust model into one with high fairness with little cost for average accuracy. We believe our research provides a meaningful contribution to the discourse on robustness and fairness in machine learning, deepening our insight into the model’s behaviors under adversarial settings.
@@ Acknowledgement
The research is supported by the UK EPSRC under project EnnCORE [EP/T026995/1].

### v2

@@ 4.2 Training from Scratch for Enhancing Fairness
Previous sections demonstrate the effectiveness and efficiency of the proposed approaches. Here we also investigate the advancements by training the model from the ground up using our method. We compare our methods with two common adversarial training methods (PGD-AT [29] and TRADES [47]) and two recent state-of-the-art techniques: CFA [40] and WAT [26], which have been proposed to mitigate the robust fairness issues recently. We adversarially trained Preact-ResNet-18 models [17] for 200 epochs with a learning rate of 0.1, which will be decayed by a factor of 0.1 at 100 and 150 epochs, successively. We start to facilitate the proposed intermediate maximization (see Algorithm 1 lines 9-14) after the 100-th epoch with the only hyper-parameter from 0.25 and enlarge it to 0.5 after the 150-th epoch. In addition, similar to CFA using weight average, we also applied EMA [21, 38], to gain a more stable performance, however, we only applied it after the 100-th epoch, where we start to apply the intermediate maximization. We report the best results under AutoAttack on the average accuracy and worst-class accuracy in Tab. 2. Besides, Fig. 3 visualizes the results of different training approaches including the proposed FAALTRADES with other 3 TRADES-based models: TRADES, CFATRADES and WATTRADES respectively. We can observe that FAAL outperforms other approaches on the worst-class clean/robust accuracy, with less sacrifice on the average robustness.
@@
| + Fine-tune with FRL-RWRM0.05 | 52.55 (22.00) | 21.11 (2.00) |
| + Fine-tune with FAALAT | 58.50 (21.00) | 21.91 (2.00) |
| + Fine-tune with FAALAT-AWP | 58.41 (19.00) | 23.44 (2.00) |
| + Fine-tune with FAALTRADES | 54.96 (18.00) | 22.71 (2.00) |
| + Fine-tune with FAALTRADES-AWP | 54.90 (18.00) | 23.25 (2.00) |
| CFATRADES | 55.57 (23.00) | 24.56 (2.00) |
| WATTRADES | 53.99 (19.00) | 22.89 (3.00) |
| FAALAT | 56.84 (16.00) | 21.85 (3.00) |
| FAALTRADES | 55.87 (21.00) | 23.57 (3.00) |
@@ 4.3 Additional Experiments on CIFAR-100 dataset
The experiments above mainly focused on CIFAR-10 dataset, which only has 10 classes in the dataset. In this section, we explore the proposed FAAL into a more challenging dataset, i.e. CIFAR-100 with 100 categories. Similarly, we reported the results of the average/worst clean accuracy and AutoAttack accuracy. The value of in our method is set to 0.05. For fine-tuning, we compare our proposed method FAAL with FRL-RWRM0.05, it can be seen that FAAL is able to achieve comparable to FRL-RWRM while reducing the amount of learning epoch up to 40 times (2 epochs vs. 80 epochs). For full adversarial training, following the experimental settings in WAT [26], we train the ResNet-18 models for 100 epochs via different adversarial training approaches, where the learning rate is decayed from 0.1 to 0.01 and 0.001 at the 75-th epoch and the 90-th epoch, respectively. We compare the results of FAAL compared to three baselines, i.e. TRADES, CFA and WAT. It can be seen in Tab. 3 that FAATTRADES achieves the highest worst-class robust accuracy (same as WAT), but it remains comparable results on the average robustness without sacrificing the average/worst-class clean accuracy. More details of the training settings can be found in the Appendix.
@@ 5 Essential Differences to SOTAs
Different form those methods that rely on historical performance or an extra validation set for manual or heuristic weight adjustment per class in each epoch, our method bypasses these requirements. FAAL introduces an additional conic convex optimization problem after the adversarial example generation, based solely on the current batch’s objective loss, the bringing solving cost is negligible. The comparison of training computation time and other key properties is illustrated in Tab. 4. As model training can be unpredictable due to random mini-batch sampling, causing quick shifts in class distribution and bias that may differ from previous epochs or validations. More importantly, FAAL can generalize to any adversarial training methods, as our intermediate maximization is a completely independent component plugged into the popular min-max framework, so it is not limited to any adversaries, unlike some methods FRL and WAT that are restricted to TRADES variants. Our data-driven component enhances flexibility in managing the balance between average robustness and robust fairness during adversarial training, and demonstrates its potential in handling various distribution shifts for the current batch.
@@ 6 Conclusion
In conclusion, we establish a connection between robust fairness and potential overfitting issues caused by the unknown group distribution shift, and present a new fairness-aware adversarial learning paradigm to address robust fairness via distributional robust optimization. Compared to state-of-the-art methods, extensive experiments on CIFAR-10 and CIFAR-100 datasets demonstrate the effectiveness and superior efficiency of the proposed approach. Notably, by just two epochs of fine-tuning, our training strategy can transform a biased robust model into one with high fairness with little cost on average accuracy. We believe our research provides a meaningful contribution to the discourse on robustness and fairness in machine learning, deepening our insight into the model’s behaviors under adversarial settings.
@@ Acknowledgement
The research is supported by the UK EPSRC under project EnnCORE [EP/T026995/1] and University of Liverpool.